OpenStudy (zacharydoc):

Solve for x: −3|x − 3| = −6 x = 1, x = −1 x = 1, x = 5 x = 0, x = 5 No solutions

OpenStudy (phi):

First, divide both sides by -3 what do you get ?

OpenStudy (zacharydoc):

1

OpenStudy (phi):

like this \[ \frac{-3\ | x-3 |}{-3}= \frac{-6}{-3} \]

OpenStudy (zacharydoc):

0 5

OpenStudy (phi):

what is -6 / -3 ?

OpenStudy (zacharydoc):

2

OpenStudy (zacharydoc):

right

OpenStudy (phi):

and what is -3 / -3 ? I assume you know that is 1 so the problem is \[ \frac{-3\ | x-3 |}{-3}= \frac{-6}{-3} \\ \text{becomes }\\ | x-3| =2 \]

OpenStudy (phi):

to solve this you assume either x-3 is positive in which case you solve x-3= 2 can you solve for x ? (I would add +3 to both sides)

OpenStudy (zacharydoc):

Ix=1,x=5

OpenStudy (phi):

then assume x-3 is negative , in which case we solve -(x-3) = 2 -x +3 = 2 -x = -1 x=1

OpenStudy (zacharydoc):

ok so b